OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x - 2)<=2? A number line with a closed circle on 3 and shading to the right. A number line with a closed circle on 3 and shading to the left. 3=>x x <= 3 help please

OpenStudy (anonymous):

can you draw it

OpenStudy (anonymous):

A, because number with 'circle on 3 and shading to the right' implies x starts at 3 and goes bigger. But \[ 2(x - 2) \le 2 \\ 2x - 4 \le 2 \\ 2x \le 6 \\ x \le 3 \] So.. X can't be bigger than 3.

OpenStudy (anonymous):

so its 3=>x

OpenStudy (anonymous):

nevermind i got it thanks